# Good practices for evaluation of synthesized speech

### v1

@@
Erica Cooper, Sébastien Le Maguer,
Esther Klabbers, Junichi Yamagishi
(March 2025)
@@
A reviewing checklist, which also serves as a summary of this document, appears on the final page.
@@ 2 Listening test evaluations in papers
It is important for papers to include some basic information about any listening tests that were conducted. This is just as important as information about datasets, training procedures, model hyperparameters, etc. and should not be omitted. The following is a list of basic information about listening tests that is necessary to report:
@@
  • Since the evaluation of the naturalness of speech is affected by differences in the volume of speech, it is necessary to be careful about the volume differences between the systems being compared. (One popular tool for volume normalization of audio samples is the sv56demo333github.com/openitu/STL/tree/dev/src/sv56 implementation of ITU P.56 [10].)

### v2

@@
Erica Cooper, Sébastien Le Maguer, Esther Klabbers, Junichi Yamagishi
(October 2025)
@@
A reviewing checklist W, also serving as a summary of this document, appears on the final page.
@@ 2 Listening test evaluations in papers
It is important for papers to include some basic information about any listening tests that were conducted. This is just as important as information about datasets, training procedures, model hyperparameters, etc. and should not be omitted. Based on [10, 11, 12], the following is a list of basic information about listening tests that is necessary to report:
@@
  • Since the evaluation of the naturalness of speech is affected by differences in the volume of speech, it is necessary to be careful about the volume differences between the systems being compared. (One popular tool for volume normalization of audio samples is the sv56demo [^3] https://github.com/openitu/STL/tree/dev/src/sv56 implementation of ITU P.56 [13].)